Name the three major shape representations and give an example for each one.


This is a Computer Graphic question.
Answer #1

1. Rectangle
This shape is defined with 2 points for height and width each.

rectangle(x1, y1, x2, y2)

The parameters defined for vertical and horizontal basis.

Example:
xVal = min( x1, x2 )
yVal = min( y1, y2 )
widthVal = abs( x1 - x2 )
heightVal = abs( y1 - y2 )
DrawRectangleFn( xVal, yVal, widthVal, heightVal)

2. Oval
It has 2 radius points of same length. It is a type of circle. It is defined and shape provided based on the angle. It can be represented as with the given formula.

(xVal + radius1 * cos(angleVal), yVal + radius2 * sin(angleVal))

3. Polygon
This shape does not have any fixed number of parameters or coordinates. It is used to define very complex shapes and structures. 2D and 3D can be represented with this shape.